The area of a rectangle is x^2-6x+58. Find x if the area of the rectangle is 50 square units.
Answer by usyim88hk(158)

x^2-6x+58 = 50 (because they are both the area of the rectangle)
-50 -50 (subtract 50 from both sides)
x^2-6x+8 = 0 (Factor it!!!)
(x-2)(x-4)= 0 (Find x!!!)
x = 2 or 4
-----------------------------------------------------------------------
Check:
if x=2
2^2-6(2)+58 = 50
4-12+58 = 50
50 = 50 (Correct!)
if x=4
4^2-6(4)+58 = 50
8-16+58 = 50
50 = 50 (Correct!)